# Learned Finite-Time Consensus for Distributed Optimization

### v1

@@ I Introduction and Related Work
Decentralized methods of the diffusion [6, 5] or consensus [7] type converge to the optimal solution under diminishing step-sizes, at the cost of convergence rate. When employing a constant step-size and exact gradients, they exhibit a non-negligible bias. This bias can be corrected with algorithms such as ADMM [8], EXTRA [9], Exact Diffusion [10], gradient tracking (GT) [11] or AugDGM [12]. The latter two correct the steady-state bias using an auxiliary variable that tracks the average gradient. Many variants of each of these algorithms exist, based on whether the graph is directed, whether the network is time-varying, and whether the communication is asynchronous [13, 14, 15]. As a representative example, the AugDGM recursion [12] is shown in (2) and (3), using the local optimization variable $w_{k,i}$, auxiliary gradient tracking variable $g_{k,i}$, step-size $\mu>0$, and combination weights at iteration $i$, $a_{k\ell}^{(i)}=[A^{(i)}]_{k,\ell}$:
@@
In [19], the convergence of linear averaging is improved by optimizing a static combination matrix $\mathcal{A}^{(i)}\triangleq\mathcal{A}$. to have the fastest mixing rate. For symmetric matrices, this is equivalent to minimizing the spectral norm, $\lVert A-\frac{1}{K}\mathbbm{1}\mathbbm{1}^{\mathsf{T}}\rVert_{2}$, which can be solved with a semi-definite program, subject to appropriate constraints on $A$. Results show improved convergence over traditional weighting rules [19].
@@
The discussion thus far motivates the use of time-varying combination matrices to improve convergence. These combination weights require specific network topologies that are known beforehand. This is also the case with the optimal learned weights in [19] and in [25], which presented algorithms for learning FT consensus weights. Furthermore, these algorithms are centralized. Motivated by these limitations, we develop a decentralized algorithm for learning optimal combination sequences for linear averaging, which do not require prior knowledge of the network topology. Numerical demonstrate that the learned combination sequences lead to improved decentralized optimization even when they are found only approximately.
@@ II Algorithm Development
We pursue the minimization of (8) using a projected gradient descent (PGD) algorithm. At each iteration, for each matrix $\{A_{j}\}_{j=1}^{\tau}$, we take a gradient step to obtain an intermediate estimate $A^{\prime(i)}_{j}$:
@@
A decentralized closed-form expression for the projection onto the intersection $\mathcal{C}_{\mathrm{sym}}\cap\mathcal{C}_{\mathrm{spars}}\cap\mathcal{C}_{\mathrm{sym}}$ does not exist. Instead, we replace the projection step in (13) with sequential projections onto so that:
@@
Having learned a sequence of combination matrices $A_{j}^{(T)}$ after $T$ iterations of the LFTC algorithm, we can then incorporate these into an averaging scheme of the form (6) by setting:
${A}^{(i)}\triangleq A_{i\%j+1}^{(T)}$ (19)
@@ III Experimental Results
We demonstrate the benefit of the sequnces (19) obtained via the LFTC Algorithm 1 in for two examples of the averaging recursion (6), as well as in the AugDGM recursions (4)–(5). The LFTC-based algorithms are compared against a static combination matrix, given by the fastest mixing rate in [19], as well as exact finite-time consensus sequences generated with full knowledge of the graph. There are a number of techniques for generating an exact finite-time consensus sequence for a given graph [24, 23, 22]. We use as a baseline the construction ${A_{s}=\frac{1}{1-\lambda_{s}}(W-\lambda_{s}I_{K})}$ where $\lambda_{s}$ are the distinct eigenvalues of $W$ less than one, and $W$ is a doubly-stochastic combination matrix consistent with the graph topology. It bears emphasising that both the fastest mixing static combination rule, and the FT consensus sequence require full central knowledge of the graph topology, and are merely used as benchmarks.
Figure 1: Linear averaging over a hypercube with 64 agents
@@
Fig. 1 shows the convergence of the averaging recursion (6) for a hypercube with 64 nodes. The proposed scheme is shown twice: after running the LFTC algorithm until convergence, and following early stopping after $100$ iterations to illustrate whether improvements are still observed when (8) is solved only approximately. The eigenvalue rule and the full LFTC algorithm have identical performance, and converge after only $6$ iterations, which corresponds to the consensus number of the hypercube. Clearly, the LFTC descent algorithm is able to learn the FT consensus sequence without any prior knowledge on the structure of the graph. The performance of the $100$-iteration LFTC algorithm does degrade, but still shows significant improvement over the static fastest mixing rate combination rule.
@@
Figure 2: Linear averaging over a Barabasi-Albert model with 32 agents
@@
Results when averaging over a 32-node Barabasi-Albert model [28] are shown in Figure 2. Compared to the hypercube, a Barabasi-Albert graph has significantly less symmetry, and a stronger community structure. The graph has a radius of 3 and so we choose a sequence length $\tau$ of 6 matrices for the LFTC. We note that compared to the hypercube, the LFTC algorithm converges more slowly for a Barabasi-Albert graph, resulting in an error of $0.1\%$ after $10^{6}$ iterations. This prevents the resulting averaging recursion (6) from converging after $6$ iterations. Nevertheless, we observe significant speedup compared to a static combination rule. The $100$-iteration LFTC is slower but still outperforms static averaging. The eigenvalue rule results in significantly less stable behavior. We attribute this to a combination of numerical errors in the eigendecomposition, resulting in inexact computation of the consensus sequence, coupled with the fact that each individual $A_{s}$ in the sequence is not necessarily constrained to being stable.
@@
Performance for a linear regression problem with a least squares objective $J_{k}(w)=\frac{1}{2}\sum_{n=1}^{N}(\gamma_{k,n}-h_{k,n}^{\mathsf{T}}w)^{2}$ on a 64-node hypercube is shown in Fig. 3. Data is generated using a linear model $\boldsymbol{\gamma}_{k}=\boldsymbol{h}_{k}^{\mathsf{T}}w^{o}+\boldsymbol{v}_{k}$ where the features $\boldsymbol{h}_{k}$follow a normal distribution $\mathcal{N}(0,I_{M})$ and $M=20$. The noise is also normally distributed with $\sigma_{v}^{2}=0.1$. Optimization follows the AugDGM algorithm in (4) and (5). Once again, the eigenvalue rule and fully-converged LFTC algorithm have identical performance, converging after 350 iterations. The $100$-iterate LFTC algorithm converges in 500 iterations, which is still significantly faster than the static combination policy.

### v2

@@ I Introduction and Related Work
Decentralized methods of the diffusion [6, 5] or consensus [7] type converge to the optimal solution under diminishing step-sizes, at the cost of convergence rate. When employing a constant step-size and exact gradients, they exhibit a non-negligible bias. This bias can be corrected with algorithms such as ADMM [8], EXTRA [9], Exact Diffusion [10], gradient tracking (GT) [11] or AugDGM [12]. The latter two correct the steady-state bias using an auxiliary variable that tracks the average gradient. Many variants of each of these algorithms exist, based on whether the graph is directed, whether the network is time-varying, and whether the communication is asynchronous [13, 14, 15]. As a representative example, the AugDGM recursion [12] is shown in (2) and (3), using the local optimization variable $w_{k}^{(i)}$, auxiliary gradient tracking variable $g_{k,i}$, step-size $\mu>0$, and combination weights at iteration $i$, $a_{k\ell}^{(i)}=[A^{(i)}]_{k,\ell}$:
@@
In [19], the convergence of linear averaging is improved by optimizing a static combination matrix $\mathcal{A}^{(i)}\triangleq\mathcal{A}$ to have the fastest mixing rate. For symmetric matrices, this is equivalent to minimizing the spectral norm, $\lVert A-\frac{1}{K}\mathbbm{1}\mathbbm{1}^{\mathsf{T}}\rVert_{2}$, which can be solved with a semi-definite program, subject to appropriate constraints on $A$. Results show improved convergence over traditional weighting rules [19].
@@
The discussion thus far motivates the use of time-varying combination matrices to improve convergence of decentralized optimization algorithms. These combination weights require specific network topologies that are known beforehand. Other works like, [19] and in [25], are centralized in their implementation. Motivated by these limitations, we develop a decentralized algorithm for learning optimal combination sequences for linear averaging, which do not require prior knowledge of the network topology. Numerical demonstrate that the learned combination sequences lead to improved decentralized optimization even when they are found only approximately.
@@ II Algorithm Development
We pursue the minimization of (8) using a projected gradient descent algorithm. At each iteration, for each matrix $\{A_{j}\}_{j=1}^{\tau}$, we take a gradient step to obtain an intermediate estimate $A^{\prime(i)}_{j}$:
@@
A decentralized closed-form expression for the projection onto the intersection $\mathcal{C}_{\mathrm{sym}}\cap\mathcal{C}_{\mathrm{spars}}\cap\mathcal{C}_{\mathrm{sym}}$ does not exist. Instead, we replace the projection step in (13) with sequential projections so that:
@@
Having learned the sequence of $\tau$ combination matrices in the LFTC algorithm, we can then incorporate these into an averaging scheme of the form (6) by setting:
$A^{(i)}\triangleq A_{i\%\tau+1}^{\mathrm{LFTC}}$ (19)
@@ III Experimental Results
We demonstrate the benefit of the sequences (19) obtained via the LFTC Algorithm 1 for two examples of the averaging recursion (6), as well as in the AugDGM recursions (4)–(5). The LFTC-based algorithms are compared against a static combination matrix, given by the fastest mixing rate in [19], as well as exact finite-time consensus sequences generated with full knowledge of the graph. There are a number of techniques for generating an exact finite-time consensus sequence for a given graph [24, 23, 22]. We use as a baseline the construction ${A_{s}=\frac{1}{1-\lambda_{s}}(W-\lambda_{s}I_{K})}$ where $\lambda_{s}$ are the distinct eigenvalues of $W$ less than one, and $W$ is a doubly-stochastic combination matrix consistent with the graph topology. It bears emphasising that both the fastest mixing static combination rule, and the FT consensus sequence require full central knowledge of the graph topology, and are merely used as benchmarks. Machine precision for the results in Fig. 1 is $10^{-15}$.
Figure 1: Linear averaging over a hypercube with 64 agents.
@@
Fig. 1 shows the mean absolute error (MAE) for the linear averaging recursion (6) for a hypercube with 64 nodes. The proposed scheme is shown twice: after running the LFTC algorithm until convergence ($150\cdot 10^{3}$ iterations), and following early stopping after $100$ iterations to illustrate whether improvements are still observed when (8) is solved only approximately. The eigenvalue rule and the full LFTC algorithm have identical performance for the linear averaging, and converge after only $6$ iterations, which corresponds to the consensus number of the hypercube. Clearly, the LFTC descent algorithm is able to learn the FT consensus sequence without any prior knowledge of the structure of the graph. The performance of the $100$-iteration LFTC algorithm does degrade, but still shows significant improvement over the static fastest mixing rate combination rule.
@@
Figure 2: Linear averaging over a Barabasi-Albert model with 32 agents.
@@
Results when averaging over a 32-node Barabasi-Albert model [28] are shown in Fig. 2. Compared to the hypercube, a Barabasi-Albert graph has significantly less symmetry, and a stronger community structure. The graph has a radius of 3 and so we choose a sequence length $\tau$ of 6 matrices for the LFTC. We note that compared to the hypercube, the LFTC algorithm converges more slowly to the optimal learned weights for a Barabasi-Albert graph, resulting in an error of $0.1\%$ between the matrix product $A_{\tau}\dots A_{2}A_{1}$ and the scaled all-ones matrix after $10^{6}$ iterations. This prevents the resulting averaging recursion (6) from converging after $6$ iterations. Nevertheless in the linear averaging recursion, we observe significant speedup compared to a static combination rule. The $100$-iteration LFTC is slower but still outperforms static averaging. The eigenvalue rule results in significantly less stable behavior. We attribute this to a combination of numerical errors in the eigendecomposition, resulting in inexact computation of the consensus sequence, coupled with the fact that each individual $A_{s}$ in the sequence is not necessarily constrained to being stable.
@@
Performance for a linear regression problem with a least squares objective $J_{k}(w)=\frac{1}{2N}\sum_{n=1}^{N}(\gamma_{k,n}-h_{k,n}^{\mathsf{T}}w)^{2}$ on a 64-node hypercube is shown in Fig. 3. Data is generated using a linear model $\boldsymbol{\gamma}_{k}=\boldsymbol{h}_{k}^{\mathsf{T}}w^{o}+\boldsymbol{v}_{k}$ where the features $\boldsymbol{h}_{k}$ follow a normal distribution $\mathcal{N}(0,I_{M})$ and $M=20$. The noise is also normally distributed with $\sigma_{v}^{2}=0.1$. Optimization follows the AugDGM algorithm in (4) and (5). Once again, the eigenvalue rule and fully-converged LFTC algorithm have identical performance, converging after 350 iterations. The $100$-iterate LFTC algorithm converges in 500 iterations, which is still significantly faster than the static combination policy.